# Retroreflective surface optimisation for optical cavities with custom mirror profiles

### v1

@@ VI Conclusion
We applied this method to optimise cavities to couple one or many emitters to the cavity mode. We showed that an optimised surface can significantly improve coupling to an emitter at the centre of a cavity by using the available mirror space more efficiently. For cases involving multiple emitters along the axis of a cavity, our method found modes with the ability to focus modes on multiple emitters simultaneously, increasing the performance in certain geometric regimes by more than an order of magnitude.
@@
Our method may find use for investigating the limits to cavity performance in alternative scenarios to those studied here, and even applications of cavities outside of cavity quantum electrodynamics. The method itself relies only upon the paraxial approximation and the assumption of low clipping loss, common operating conditions for Fabry-Perot cavities, offering considerable flexibility over which metric is optimised.
VII Acknowledgements
This work was funded by the UK Engineering and Physical Sciences Research Council Hub in Quantum Computing and Simulation (EP/T001062/1) and Hub for Quantum Computing via Integrated and Interconnected Implementations (EP/Z53318X/1). The authors acknowledge the use of the IRIDIS High Performance Computing Facility, and associated support services at the University of Southampton, in the completion of this work. Data underlying the results presented in this paper are available at DOI added on acceptance. The authors declare no conflicts of interest.
Appendix A Gradient-based retroreflective optimisation
@@
$m=p\frac{\Pi_{i=1}^{i_{M}}\vec{C}^{T}\hat{N}^{(i)}\vec{C}}{\Pi_{j=1}^{j_{M}}\vec{C}^{T}\hat{D}^{(j)}\vec{C}}\frac{\left(\vec{C}^{T}\vec{C}\right)^{j}}{\left(\vec{C}^{T}\vec{C}\right)^{i}},$ (17)
where $\hat{N}^{(i)}$ is the $i$-th matrix of $i_{M}$ whose expectation is on the numerator, and $\hat{D}^{(j)}$ is the $j$-th matrix of $j_{M}$ whose expectation is on the denominator, and $p$ is a scalar prefactor. The factor on the right hand side is a normalisation so that the metric is independent of the overall scale of the mode coefficients.
For the cases in this manuscript, we assume that $\epsilon_{\mathrm{T}}$ is symmetric upon an axial flip (i.e. has the same intensity at two points related by a reflection in the $xy$-plane through the origin). This is done so that the two mirrors of the cavity are mirror images of each other (and in our case, identical due to each mirror’s azimuthal rotation symmetry), which considerably simplifies the analysis. In this case, the coefficients of $\vec{C}$ are all real, and we will derive expressions according to this case, but one could equally perform the derivation without this assumption.
The two derivatives we require are
$\vec{\nabla}\left(\vec{C}^{T}\vec{C}\right)=2\vec{C},\quad\vec{\nabla}\left(\vec{C}^{T}\hat{M}\vec{C}\right)=2\mathcal{R}(\hat{M}\vec{C}),$ (18)
where $\mathcal{R}$ signifies the element-wise real part.
We can then use the generic formula
@@
This gradient is supplied to the optimisation routine for many of our example cases.
Appendix B Justification of the use of retroreflective optimisation
@@
Appendix C Derivation of matrices for the optimising photon extraction
As discussed in Sec. III of the main text, the retroreflective optimisation can proceed more smoothly if quantities within the evaluated metric can be calculated as expectation values of Hermitian matrices. For the case of photon extraction, our chosen metric is the internal cooperativity $C_{\mathrm{int}}$, which has expression (taken from Eq. 10 of the main text, but repeated here)
where $\lambda$ is the wavelength, $\mathcal{L}_{\mathrm{int}}$ is the fractional round trip loss to non-transmissive sources, $\bm{r_{e}}$ is the position of the emitter, and $A^{-1}_{\mathrm{eff}}(\bm{r})$ is the normalised intensity of the forward running wave at position $\bm{r}$,
$A^{-1}_{\mathrm{eff}}(\bm{r})=\frac{I(\bm{r})}{\int_{\Pi_{z_{r}}}I(\bm{r^{\prime}})\,d^{2}\bm{r^{\prime}}},$ (22)
where $\Pi_{z_{r}}$ is the transverse ($xy$) plane that passes through $\bm{r}$ and $I(\bm{r})$ is the intensity of the electric field at position $\bm{r}$. Of the quantities in Eq. 21, the inverse effective area $A^{-1}_{\mathrm{eff}}(\bm{r_{e}})$, and internal loss $\mathcal{L}_{\mathrm{int}}$ depend upon the cavity mode, and we therefore need to find matrices $\hat{A}^{-1}_{\mathrm{eff.}}(\bm{r_{e}})$ and $\hat{\mathcal{L}}_{\mathrm{int}}$ such that
$A^{-1}_{\mathrm{eff}}(\bm{r_{e}})=\frac{\vec{C}^{T}\hat{A}^{-1}_{\mathrm{eff.}}(\bm{r_{e}})\vec{C}}{\vec{C}^{T}\vec{C}},\quad\mathcal{L}_{\mathrm{int}}=\frac{\vec{C}^{T}\hat{\mathcal{L}}_{\mathrm{int}}\vec{C}}{\vec{C}^{T}\vec{C}},$ (23)
where the normalisation factor is an acknowledgment that the quantities $A^{-1}_{\mathrm{eff}}(\bm{r})$ and $\mathcal{L}_{\mathrm{int}}$ are (by design) independent of the total intensity of the mode.
The first example is the inverse effective area matrix. Using the expression for electric field intensity
@@ Appendix C Derivation of matrices for the optimising photon extraction
where $E(\bm{r})$ is the temporally constant electric field magnitude at position $\bm{r}$, and Eqs. 1 and 5 from the main text, yields
$\hat{A}^{-1}_{\mathrm{eff.}}(\bm{r})$ $=\frac{\sum_{s^{\prime}=0}^{S}\sum_{s=0}^{S}\left(C_{s^{\prime}}\right)^{*}\left(\epsilon^{+}_{s^{\prime}}(\bm{r})\right)^{*}\epsilon^{+}_{s}(\bm{r})C_{s}}{\int_{\Pi_{z_{r}}}\sum_{s^{\prime}=0}^{S}\sum_{s=0}^{S}\left(C_{s^{\prime}}\right)^{*}\left(\epsilon^{+}_{s^{\prime}}(\bm{r^{\prime}})\right)^{*}\epsilon^{+}_{s}(\bm{r^{\prime}})C_{s}\,d^{2}\bm{r^{\prime}}},$ (25)
@@
$\left(\hat{A}^{-1}_{\mathrm{eff.}}(\bm{r})\right)_{s^{\prime},s}$ $=\left(\epsilon^{+}_{s^{\prime}}(\bm{r})\right)^{*}\epsilon^{+}_{s}(\bm{r}),$
@@
where the integrals in the denominator of the first line are simplified using the knowledge that $\{\epsilon_{s}\}$ is an orthonormal basis in every transverse plane.
To calculate the internal cooperativity, an evaluation of the internal loss $\mathcal{L}_{\mathrm{int}}$ in also required. To understand how this is calculated, suppose a running wave field $\epsilon^{+}$ propagates towards a mirror at axial coordinate $z_{m}$ with a finite surface $\Pi_{m}$ and combined scattering and absorption losses $\mathcal{L}_{\mathrm{int}}^{\mathrm{clip}\rightarrow 0}/2$, to match the definition of $\mathcal{L}_{\mathrm{int}}^{\mathrm{clip}\rightarrow 0}$ from the main text as the round trip non-clipping internal loss. The power losses sustained upon this reflection will be identical to those sustained propagating through an aperture with the same domain as the mirror surface, where the field outside the aperture is completely blocked, and the field inside the aperture incurs an intensity loss $\mathcal{L}_{\mathrm{int}}^{\mathrm{clip}\rightarrow 0}/2$. We can therefore write the fractional internal intensity loss caused by the mirror from the field propagation in this analogous scenario through
@@
$\frac{1}{2}\hat{\mathcal{L}}_{\mathrm{int}}=\frac{1-\int_{\Pi_{z_{m}+\epsilon}}\left(\epsilon^{+}(\bm{r})\right)^{*}\epsilon^{+}(\bm{r})\,d\Pi}{\int_{\Pi_{z_{m}-\epsilon}}\left(\epsilon^{+}(\bm{r})\right)^{*}\epsilon^{+}(\bm{r})\,d\Pi},$ (26)
where $\Pi_{z_{m}-\epsilon}$ is the infinite transverse plane just before the mirror, $\Pi_{z_{m}+\epsilon}$ the infinite plane after the mirror aperture, and the factor of $1/2$ before $\mathcal{L}_{\mathrm{int}}$ signifies that the loss fraction from one mirror is half the round trip loss. The coefficients $\vec{C}^{\Pi_{z_{m}+\epsilon}}$ of the mode superposition after the mirror can be written in terms of those of the cavity mode ($\vec{C}$) with
$\vec{C}^{\Pi_{z_{m}+\epsilon}}_{s^{\prime}}$ $=\sqrt{1-\frac{1}{2}\mathcal{L}_{\mathrm{int}}^{\mathrm{clip}\rightarrow 0}}\sum_{s=0}^{S}\hat{\rho}_{s^{\prime},s}\vec{C}_{s},$ (27)
$\hat{\rho}_{s^{\prime},s}$ $=\int_{\Pi_{m}}\left(\epsilon^{+}_{s^{\prime}}(\bm{r})\right)^{*}\epsilon^{+}_{s}(\bm{r})\,d^{2}\bm{r},$ (28)
where $\hat{\rho}$ is the clipping matrix of the aperture defined by the mirror edge. For the spherical apertures considered in this manuscript, the elements of $\hat{\rho}$ can be evaluated using the techniques of Vajente [55], where elevated-precision numerical evaluation is required. Combining Eq. 26 with Eq. 5 from the main text, and substituting Eq. 27, leaves the fractional non-transmissive loss induced by the mirror as
@@
$\hat{\mathcal{L}}_{\mathrm{int}}=2\frac{\vec{C}^{T}\left[\hat{\mathcal{I}}-\left(1-\frac{1}{2}\mathcal{L}_{\mathrm{int}}^{\mathrm{clip}\rightarrow 0}\right)\hat{\rho}^{{\dagger}}\hat{\rho}\right]\vec{C}}{\vec{C}^{T}\vec{C}},$ (29)
where $\hat{\mathcal{I}}$ is the identity matrix, which is, again, a normalised expectation value over a Hermitian matrix.
@@
Appendix D Derivation of critical diameter, separation, and best geometry for coupling to two emitters

### v2

@@ VII Conclusion
We applied this method to optimise cavities to couple one or many emitters to the cavity mode. We showed that an optimised surface can significantly improve coupling to a single emitter, either central in the cavity or displaced axially or transversely, by using the available mirror space more efficiently. For cases involving multiple emitters along the axis of a cavity, our method found modes with the ability to focus modes on multiple emitters simultaneously, increasing the performance in certain geometric regimes by more than an order of magnitude.
@@
Finally, our method itself relies only upon the paraxial approximation and the assumption of low clipping loss, common operating conditions for Fabry-Perot cavities. This offers considerable flexibility over which metric is optimised, and could allow for the optimisation of cavity geometries in alternative scenarios to those studied here, such as collective coupling to probe ultrastrong coupling [75] or many-body physics [76, 77], using vacuum coupling to control chemical reactions [78, 79, 80], or even applications outside cavity quantum electrodynamics.
Acknowledgements.
This work was funded by the UK Engineering and Physical Sciences Research Council Hub in Quantum Computing and Simulation (EP/T001062/1) and Hub for Quantum Computing via Integrated and Interconnected Implementations (EP/Z53318X/1). The authors acknowledge the use of the IRIDIS High Performance Computing Facility, and associated support services at the University of Southampton, in the completion of this work. Data underlying the results presented in this paper are available at Ref. [81]. The authors declare no conflicts of interest.
Appendix A Role of internal cooperativity in photon extraction
Though the retroreflective optimisation method described in the main text can optimise cavity modes for arbitrary figures of merit, the main text focuses on optimisation of the internal cooperativity due to its critical role in many protocols. Specifically for single photon generation over adiabatic timescales, $C_{\mathrm{int}}$ is the single figure of merit that limits the extraction probability
$P_{\mathrm{ext}}=\frac{\sqrt{2C_{\mathrm{int}}+1}-1}{\sqrt{2C_{\mathrm{int}}+1}+1},$ (20)
originally written as Eq. (10) in the main text [49]. This extraction probability is visualised in Fig. 8 so that the practical impact of optimising the internal cooperativity can be understood. For internal cooperativities below 10, increases in the internal cooperativity result in strong increases in $P_{\mathrm{ext}}$ (Fig. 8 blue line). For internal cooperativities above 10, $P_{\mathrm{ext}}$ is already a significant fraction of unity, but improvements will dramatically reduce the probability that a photon is not generated (Fig. 8 red line). The benefits to improvements in the internal cooperativity thus depend upon the $C_{\mathrm{int}}$ of the system before optimisation. It should, however, be noted that the value for $C_{\mathrm{int}}$ quoted in this manuscript (Eq. (11)) assumes unit branching ratio to the final state. A real system with non-unity branching ratio will have $C_{\mathrm{int}}$ reduced by a factor of the branching ratio.
Figure 8: The adiabatic upper limit to the probability that a photon is extracted (blue), and the lower limit to the probability that a photon is not extracted (red) from an emitter-cavity system with internal cooperativity $C_{\mathrm{int}}$ whose output transmission is optimised for photon extraction probability at each datapoint.
Appendix B Gradient-based retroreflective optimisation
@@
$m=p\left[\frac{\Pi_{i=1}^{i_{M}}\vec{C}^{T}\hat{N}^{(i)}\vec{C}}{\Pi_{j=1}^{j_{M}}\vec{C}^{T}\hat{D}^{(j)}\vec{C}}\right]\frac{\left(\vec{C}^{T}\vec{C}\right)^{j_{M}}}{\left(\vec{C}^{T}\vec{C}\right)^{i_{M}}},$ (21)
where $\hat{N}^{(i)}$ is the $i$-th Hermitian matrix of $i_{M}$ whose expectation is on the numerator, and $\hat{D}^{(j)}$ is the $j$-th Hermitian matrix of $j_{M}$ whose expectation is on the denominator, and $p$ is a scalar prefactor. The factor on the right hand side is a normalisation so that the metric is independent of the overall scale of the mode coefficients.
For most of the cases in this manuscript, the geometry is symmetric under $z\rightarrow-z$, and we therefore assume that $|\absolutevalue{\epsilon_{\mathrm{T}}}$ is unchanged upon the transformation $z\rightarrow-z$. In this case, the coefficients of $\vec{C}$ all have the same phase, and thus we choose that all coefficients are real. We will first derive expressions for gradient-based optimisation under this assumption, because it is easy to then adapt those results to cases without this symmetry.
The two derivatives we require are, on the assumption that $\vec{C}$ is a real vector,
[…]
@@
Now let us consider the more general case that $\vec{C}$ has complex coefficients. We can turn this into a real vector $\vec{C}^{\mathcal{R}}$ with twice the number of coefficients by defining
[…]
Appendix C Justification of the use of retroreflective optimisation
@@
Appendix D Derivation of matrices for the optimising photon extraction
As discussed in Sec. III of the main text, the retroreflective optimisation can proceed more smoothly if quantities within the evaluated metric can be calculated as expectation values of Hermitian matrices. For the case of photon extraction, our chosen metrics are the internal cooperativity $C_{\mathrm{int}}$, and its single-transverse dimension equivalent $C_{\mathrm{int}}/L^{-1}_{\mathrm{y,eff}}$, which have expressions (taken from Eqs. (11) and (12) of the main text, but repeated here)
$C_{\mathrm{int}}=\frac{3\lambda^{2}A^{-1}_{\mathrm{eff}}(\bm{r_{e}})}{\pi\mathcal{L}_{\mathrm{int}}},\quad C_{\mathrm{int}}/L^{-1}_{\mathrm{y,eff}}=\frac{3\lambda^{2}L^{-1}_{\mathrm{eff}}(\bm{r_{e}})}{\pi\mathcal{L}_{\mathrm{int}}},$ (28)
where $\lambda$ is the wavelength, $\mathcal{L}_{\mathrm{int}}$ is the fractional round trip loss to non-transmissive sources, $\bm{r_{e}}$ is the position of the emitter, and $A^{-1}_{\mathrm{eff}}(\bm{r})$ ($L^{-1}_{\mathrm{eff}}(\bm{r})$) is the normalised intensity per area (length) of the forward running wave at position $\bm{r}$,
$A^{-1}_{\mathrm{eff}}(\bm{r})\,\,\mathrm{or}\,\,L^{-1}_{\mathrm{eff}}(\bm{r})=\frac{I(\bm{r})}{\int_{\Pi_{z_{r}}}I(\bm{r^{\prime}})\,d\Pi_{z_{r}}},$ (29)
where $\Pi_{z_{r}}$ is used to unify the notation for the two cases by representing the infinite transverse domain ($xy$-plane or $x$-axis for two or one transverse dimensions respectively) that passes through $\bm{r}$, and $I(\bm{r})$ is the intensity of the electric field at position $\bm{r}$ (per area for two transverse dimensions, and per length for a single transverse dimension). Of the quantities in Eq. (28), the inverse effective area $A^{-1}_{\mathrm{eff}}(\bm{r_{e}})$, inverse effective length $L^{-1}_{\mathrm{eff}}(\bm{r_{e}})$, and internal loss $\mathcal{L}_{\mathrm{int}}$ depend upon the cavity mode, and we therefore need to find matrices $\hat{A}^{-1}_{\mathrm{eff}}(\bm{r_{e}})$ (or $\hat{L}^{-1}_{\mathrm{eff}}(\bm{r_{e}})$ for single transverse dimension), and $\hat{\mathcal{L}}_{\mathrm{int}}$ such that
$A^{-1}_{\mathrm{eff}}(\bm{r_{e}})=\frac{\vec{C}^{T}\hat{A}^{-1}_{\mathrm{eff}}(\bm{r_{e}})\vec{C}}{\vec{C}^{T}\vec{C}},\quad L^{-1}_{\mathrm{eff}}(\bm{r_{e}})=\frac{\vec{C}^{T}\hat{L}^{-1}_{\mathrm{eff}}(\bm{r_{e}})\vec{C}}{\vec{C}^{T}\vec{C}},\quad\mathcal{L}_{\mathrm{int}}=\frac{\vec{C}^{T}\hat{\mathcal{L}}_{\mathrm{int}}\vec{C}}{\vec{C}^{T}\vec{C}},$ (30)
where the normalisation factor is an acknowledgment that the quantities $A^{-1}_{\mathrm{eff}}(\bm{r})$, $L^{-1}_{\mathrm{eff}}(\bm{r})$, and $\mathcal{L}_{\mathrm{int}}$ are (by design) independent of the total intensity of the mode.
The first example is the inverse effective area matrix (used for the radially-symmetric situations that comprise the majority of the manuscript). Using the expression for electric field intensity
@@ Appendix D Derivation of matrices for the optimising photon extraction
where $E(\bm{r})$ is the temporally constant electric field magnitude at position $\bm{r}$, and Eqs. (1) and (6) from the main text, yields
$\hat{A}^{-1}_{\mathrm{eff}}(\bm{r})$ $=\frac{\sum_{s^{\prime}=0}^{S}\sum_{s=0}^{S}\left(C_{s^{\prime}}\right)^{*}\left(\epsilon^{LG,+}_{s^{\prime}}(\bm{r})\right)^{*}\epsilon^{LG,+}_{s}(\bm{r})C_{s}}{\int_{\Pi_{z_{r}}}\sum_{s^{\prime}=0}^{S}\sum_{s=0}^{S}\left(C_{s^{\prime}}\right)^{*}\left(\epsilon^{LG,+}_{s^{\prime}}(\bm{r^{\prime}})\right)^{*}\epsilon^{LG,+}_{s}(\bm{r^{\prime}})C_{s}\,d^{2}\bm{r^{\prime}}},$ (32)
@@
$\left(\hat{A}^{-1}_{\mathrm{eff}}(\bm{r})\right)_{s^{\prime},s}$ $=\left(\epsilon^{LG,+}_{s^{\prime}}(\bm{r})\right)^{*}\epsilon^{LG,+}_{s}(\bm{r}),$
@@
where the integrals in the denominator of the first line are simplified using the knowledge that $\{\epsilon_{s}^{LG,+}\}$ is an orthonormal basis in every transverse plane.
An exactly analogous derivation holds for $\hat{L}^{-1}_{\mathrm{eff}}$ on the understanding that, consistent with Eqs. (1) and (3) in the main text and the conventions in this appendix, the dimensions of $I$ and $E$ change upon moving to a single transverse dimension. This means that we can use the more compact expression
$\left(\hat{A}^{-1}_{\mathrm{eff}}(\bm{r})\right)_{s^{\prime},s}\,\,\mathrm{or}\,\,\left(\hat{L}^{-1}_{\mathrm{eff}}(\bm{r})\right)_{s^{\prime},s}=\left(\epsilon^{\Lambda,+}_{s^{\prime}}(\bm{r})\right)^{*}\epsilon^{\Lambda,+}_{s}(\bm{r}),$ (33)
on the understanding that $\Lambda=\{LG,HG\}$ is chosen to match the transverse dimensionality of the scenario.
To calculate the internal cooperativity, an evaluation of the internal loss $\mathcal{L}_{\mathrm{int}}$ in also required. For this, it is important to clarify the bases of the propagating modes. Conventional mode mixing expresses the mirror at positive (negative) $z$ as a matrix that scatters from the positively (negatively) propagating basis to the negatively (positively) propagating basis. Instead, we want to find our matrices so that they operate purely on the mode coefficients in the positively propagating basis, with the negatively propagating mode completely specified by the assumption of perfect retroreflection (this avoids complications in the Laguerre-Gauss basis where retroreflection swaps coefficients from $m\rightarrow-m$ because retroreflection preserves helicity not angular momentum). With this in mind, it is easiest to start with the mirror at positive $z$. Imagine replacing the mirror with an aperture of the same size, where light passing through the aperture sustains combined scattering and absorption losses $\mathcal{L}_{\mathrm{abs}}/2$, to match the definition of $\mathcal{L}_{\mathrm{abs}}$ from the main text as the round trip non-clipping internal loss. The loss upon reflection from the mirror is the same as the loss of power upon this imagined transmission as the only further action of the mirror is to redirect power. The coefficients $\vec{C}_{\mathrm{out}}$ of the mode superposition representing this imagined transmission can be written in terms of the input coefficients ($\vec{C}$) with
@@
$\left(\vec{C}_{\mathrm{out}}\right)_{s^{\prime}}=\sqrt{1-\frac{1}{2}\mathcal{L}_{\mathrm{abs}}}\sum_{s=0}^{S}\hat{\rho}_{s^{\prime},s}\vec{C}_{s},\quad\hat{\rho}^{+}_{s^{\prime},s}=\int_{\Pi_{m_{+}}}\left(\epsilon^{\mathrm{\Lambda,+}}_{s^{\prime}}(\bm{r})\right)^{*}\epsilon^{\mathrm{\Lambda,+}}_{s}(\bm{r})\,d\Pi,$ (34)
where $\hat{\rho}^{+}$ is the clipping matrix of the aperture defined by the edge of the mirror at positive $z$ (which is to what its ++ superscript refers), and $\Pi_{m_{+}}$ is the finite domain of the mirror at positive $z$. The internal loss matrix of the mirror can then be derived
$\mathcal{L}_{\mathrm{int}}^{+}=\frac{\left(\vec{C}\right)^{{\dagger}}\vec{C}-\left(\vec{C}_{\mathrm{out}}^{L/R}\right)^{{\dagger}}\vec{C}_{\mathrm{out}}^{L/R}}{\left(\vec{C}\right)^{{\dagger}}\vec{C}},\quad\hat{\mathcal{L}}_{\mathrm{int}}^{+}=\hat{\mathcal{I}}-\left(1-\frac{1}{2}\mathcal{L}_{\mathrm{abs}}\right)\hat{\rho^{+}}^{{\dagger}}\hat{\rho^{+}}$, (35)
For the mirror at negative $z$, we can repeat this process with the the negatively propagating mode $\epsilon^{-}$ to see that we would have
@@
$\left[\hat{\mathcal{L}}_{\mathrm{int}}^{-}\right]_{\mathrm{neg.basis}}=\hat{\mathcal{I}}-\left(1-\frac{1}{2}\mathcal{L}_{\mathrm{abs}}\right)\hat{\rho^{-}}^{{\dagger}}\hat{\rho^{-}},\quad\hat{\rho}^{-}_{s^{\prime},s}=\int_{\Pi_{m_{-}}}\left(\epsilon^{\mathrm{\Lambda,-}}_{s^{\prime}}(\bm{r})\right)^{*}\epsilon^{\mathrm{\Lambda,-}}_{s}(\bm{r})\,d\Pi,$ (36)
for an expression of $\hat{\mathcal{L}}_{\mathrm{int}}^{-}$ in the negative propagating basis (whatever those basis states happen to be). To convert this to the positive propagating basis, we note that $\hat{\rho}$ is an integral of the normalised intensity matrix (i.e. inverse effective area or inverse effective length) over the surface of the mirror (see Eq. (33)). Because reflection does not translate power transversely, within the domain of the mirrors we know that the intensity of the incoming mode and outgoing mode are the same everywhere. Therefore, over the mirror domain we can replace the intensity matrix of $\epsilon^{\Lambda,-}$ with the intensity matrix of $\epsilon^{\Lambda,+}$. Combining these results, we can write the internal loss matrices of both mirrors as
$\hat{\mathcal{L}}_{\mathrm{int}}^{\pm}=\hat{\mathcal{I}}-\left(1-\frac{1}{2}\mathcal{L}_{\mathrm{abs}}\right)\hat{\rho^{\pm}}^{{\dagger}}\hat{\rho^{\pm}},\quad\hat{\rho}^{\pm}_{s^{\prime},s}=\int_{\Pi_{m_{\pm}}}\left(\epsilon^{\mathrm{\Lambda,+}}_{s^{\prime}}(\bm{r})\right)^{*}\epsilon^{\mathrm{\Lambda,+}}_{s}(\bm{r})\,d\Pi,$ (37)
which is now a form that takes mode coefficients purely in the positive propagating basis. In the case that the mirror apertures have identical diameter and distance from the origin, and that the mode is restricted to be even symmetric with respect to $z$-reflection about the origin, the internal loss matrices for the two mirrors are identical. This can help reduce calculation time.
@@
Appendix E Derivation of critical diameter, separation, and best geometry for coupling to two emitters